# You May Use the Same Channel Knowledge Map for Environment-Aware NLoS Sensing and Communication

### v2

@@ IV-B CADM-enabled sensing
$+\frac{\theta_{l}-\mu_{\theta,{l_{R}}}(\mathbf{x})}{\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})}\nabla_{\mathbf{x}}\mu_{\theta,{l_{R}}}(\mathbf{x})+\frac{\left(\theta_{l}-\mu_{\theta,{l_{R}}}(\mathbf{x})\right)^{2}-\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})}{2\sigma_{\theta,{l_{R}}}^{4}(\mathbf{x})}\nabla_{\mathbf{x}}\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})+\frac{\tau_{l}-\mu_{\tau,{l_{T}}}(\mathbf{x})-\mu_{\tau,{l_{R}}}(\mathbf{x})}{\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})}\nabla_{\mathbf{x}}\big(\mu_{\tau,{l_{T}}}(\mathbf{x})+\mu_{\tau,{l_{R}}}(\mathbf{x})\big)$
$+\frac{\left(\tau_{l}-\mu_{\theta,{l_{T}}}(\mathbf{x})-\mu_{\theta,{l_{R}}}(\mathbf{x})\right)^{2}-\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})-\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})}{2(\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x}))^{2}}\nabla_{\mathbf{x}}\big(\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})\big)\bigg).$
To estimate the sensing target location $\mathbf{x}$ using gradient descent as specified in (12), the gradient of $\ln P(\hat{\mathbf{z}}|\mathbf{x})$ with respect to $\mathbf{x}$ is required. This gradient is given by (41) shown on the top of the next page, where the gradients $\nabla_{\mathbf{x}}\mu_{\theta,l_{T}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\theta,l_{T}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\tau,l_{T}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\tau,l_{T}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\theta,l_{R}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\theta,l_{R}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\tau,l_{R}}(\mathbf{x})$, and $\nabla_{\mathbf{x}}\sigma_{\tau,l_{R}}^{2}(\mathbf{x})$ can be obtained based on the structure FCNN as (37) by substituting $\mathbf{q}$ with $\mathbf{x}$. For example, the gradient of a parameter such as $\mu_{\theta,l_{T}}(\mathbf{x})$ is written as
@@
As a summary, to estimate the location $\mathbf{x}$ of the sensing target for the challenging NLoS environment, parameters such as AoD, AoA, and delay for each path are firstly estimated based on the echo signal $\mathbf{y}(t)$ in (6) using well established algorithms (e.g., MUSIC, ESPRIT, or compressed sensing [8]), forming the observation vector $\hat{\mathbf{z}}$. The methods cited, such as MUSIC or ESPRIT, are mentioned only as well-established, high-resolution examples, not as mandatory components. For dynamic scenarios involving high-speed targets, where low latency and real-time capability are required, our framework can be seamlessly integrated with various estimation algorithms (e.g., FFT-based) specifically designed for high-mobility tracking. The core contribution of this work lies in intelligently utilizing the outputs $\hat{\mathbf{z}}$ of any such algorithms for NLoS localization. Then, substituting $\hat{\mathbf{z}}$ into (41) allows computation of the gradient, enabling estimation of the target location $\mathbf{x}$ via gradient descent in (12). Note that while the algorithm is presented for the most challenging NLoS scenarios, it is applicable for the LoS case as well. The pseudo-code of CKM-enabled environment-aware sensing is elaborated in Algorithm 1.
@@
7:  Update sensing target location estimation: $\mathbf{x}^{(k+1)}=\mathbf{x}^{(k)}-\eta\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))\big|_{\mathbf{x}^{(k)}}$
8:  $k\leftarrow k+1$
9:end while
10:Return $\hat{\mathbf{x}}=\mathbf{x}^{(k)}$
@@ V Simulation Results
To construct the CKM, we use a FCNN as shown in Fig. 6. It consists of five layers: an input layer with 2 neurons (corresponding to the 2D coordinates of UE location $\mathbf{q}$), three hidden layers with $128\times 256\times 128$ neurons using LeakyReLU activation with a negative slope of 0.01, and an output layer with $4L^{\prime}$ neurons, where $L^{\prime}=5$ is the number of Gaussian components in (36). The output includes the angle mean $\mu_{\theta,l^{\prime}}$, the angle variance $\sigma_{\theta,l^{\prime}}^{2}$, the delay mean $\mu_{\tau,l^{\prime}}$, and the delay variance $\sigma_{\tau,l^{\prime}}^{2}$ for each path $l^{\prime}$. Besides, the training dataset is generated by sampling 10,000 points uniformly within the 100 m $\times$ 100 m area. This dataset was empirically found to be sufficient for the FCNN to converge and was partitioned into a training set (80%) and a validation set (20%). For each sampled UE location in the training dataset, the ground-truth channel is generated by considering all 40 possible single-bounce paths via the scatterers. The delay for each path is calculated based on its total geometric length, which is the distance from the BS to the scatterer and then to the UE. The complex gain of each path is determined by two main factors: the path-length dependent attenuation and a unique random reflection coefficient assigned to each scatterer. Although 40 paths are generated, the CADM is designed to learn the $L^{\prime}=5$ dominant ones. Therefore, a path selection mechanism is applied: the 5 paths with the highest power, as determined by their path gains, are chosen. The angles and delays of these 5 selected paths constitute the ground-truth data used for training the FCNN and the training loss is computed as the mean squared error (MSE) between the predicted and true channel parameters. The network was trained for 500 epochs using the Adam optimizer, with an initial learning rate of 0.01 and a batch size of 128.
@@
  • NLoS, Geometry-based: Assume the shortest-delay path as LoS for geometric localization.
@@ VI Conclusion
This paper introduces a CKM-enabled ISAC framework that achieves robust environment-aware sensing, even in challenging NLoS scenarios where traditional geometric-based methods fail. We demonstrate that the same CKM designed for environment-aware communications can be exploited to achieve sensing as well. Simulation results demonstrate significant performance gains over geometry-based approaches, with CRLB analysis confirming enhanced robustness. This approach marks a significant advancement for ISAC systems, enabling seamless integration of sensing and communication in complex environments enabled by the same CKM.

### v3

@@ IV-B CADM-enabled sensing
$+\frac{\theta_{l}-\mu_{\theta,{l_{R}}}(\mathbf{x})}{\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})}\nabla_{\mathbf{x}}\mu_{\theta,{l_{R}}}(\mathbf{x})+\frac{\left(\theta_{l}-\mu_{\theta,{l_{R}}}(\mathbf{x})\right)^{2}-\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})}{2\sigma_{\theta,{l_{R}}}^{4}(\mathbf{x})}\nabla_{\mathbf{x}}\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x})$
$+\frac{\tau_{l}-\mu_{\tau,{l_{T}}}(\mathbf{x})-\mu_{\tau,{l_{R}}}(\mathbf{x})}{\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})}\nabla_{\mathbf{x}}\big(\mu_{\tau,{l_{T}}}(\mathbf{x})+\mu_{\tau,{l_{R}}}(\mathbf{x})\big)$
$+\frac{\left(\tau_{l}-\mu_{\tau,{l_{T}}}(\mathbf{x})-\mu_{\tau,{l_{R}}}(\mathbf{x})\right)^{2}-\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})-\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})}{2(\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x}))^{2}}\nabla_{\mathbf{x}}\big(\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})\big)\bigg).$
To estimate the sensing target location $\mathbf{x}$ using gradient descent as specified in (12), the gradient of $\ln P(\hat{\mathbf{z}}|\mathbf{x})$ with respect to $\mathbf{x}$ is required. This gradient is given by (41) shown on the top of the next page, where the gradients $\nabla_{\mathbf{x}}\mu_{\theta,l_{T}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\theta,l_{T}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\tau,l_{T}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\tau,l_{T}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\theta,l_{R}}(\mathbf{x})$, $\nabla_{\mathbf{x}}\sigma_{\theta,l_{R}}^{2}(\mathbf{x})$, $\nabla_{\mathbf{x}}\mu_{\tau,l_{R}}(\mathbf{x})$, and $\nabla_{\mathbf{x}}\sigma_{\tau,l_{R}}^{2}(\mathbf{x})$ can be obtained based on the structure of FCNN as (37) by substituting $\mathbf{q}$ with $\mathbf{x}$. For example, the gradient of a parameter such as $\mu_{\theta,l_{T}}(\mathbf{x})$ is written as
@@
As a summary, to estimate the location $\mathbf{x}$ of the sensing target for the challenging NLoS environment, parameters such as AoD, AoA, and delay for each path are firstly estimated based on the echo signal $\mathbf{y}(t)$ in (6) using well-established algorithms (e.g., MUSIC, ESPRIT, or compressed sensing [3]), forming the observation vector $\hat{\mathbf{z}}$. The methods cited, such as MUSIC or ESPRIT, are mentioned only as well-established, high-resolution examples, not as mandatory components. For dynamic scenarios involving high-speed targets, where low latency and real-time capability are required, our framework can be seamlessly integrated with various estimation algorithms (e.g., FFT-based) specifically designed for high-mobility tracking. The core contribution of this work lies in intelligently utilizing the outputs $\hat{\mathbf{z}}$ of any such algorithms for NLoS localization. Then, substituting $\hat{\mathbf{z}}$ into (41) allows computation of the gradient, enabling estimation of the target location $\mathbf{x}$ via gradient descent in (12). Note that while the algorithm is presented for the most challenging NLoS scenarios, it is applicable for the LoS case as well. The pseudo-code of CKM-enabled environment-aware sensing is elaborated in Algorithm 1.
@@
7:  Update sensing target location estimation:
8:   $\mathbf{x}^{(k+1)}=\mathbf{x}^{(k)}-\eta\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))\big|_{\mathbf{x}^{(k)}}$
9:  $k\leftarrow k+1$
10:end while
11:Return $\hat{\mathbf{x}}=\mathbf{x}^{(k)}$
@@ V Simulation Results
To construct the CKM, we use an FCNN as shown in Fig. 6. It consists of five layers: an input layer with 2 neurons (corresponding to the 2D coordinates of UE location $\mathbf{q}$), three hidden layers with $128\times 256\times 128$ neurons using LeakyReLU activation with a negative slope of 0.01, and an output layer with $4L^{\prime}$ neurons, where $L^{\prime}=5$ is the number of Gaussian components in (36). The output includes the angle mean $\mu_{\theta,l^{\prime}}$, the angle variance $\sigma_{\theta,l^{\prime}}^{2}$, the delay mean $\mu_{\tau,l^{\prime}}$, and the delay variance $\sigma_{\tau,l^{\prime}}^{2}$ for each path $l^{\prime}$. Besides, the training dataset is generated by sampling 10,000 points uniformly within the 100 m $\times$ 100 m area. This dataset was empirically found to be sufficient for the FCNN to converge and was partitioned into a training set (80%) and a validation set (20%). Note that in practical deployments, it is unnecessary to physically collect all 10,000 samples. As mentioned in Section I, advanced CKM construction methods based on generative AI [8, 12] can complete a large-scale dataset from only a small fraction of real-world measurements, significantly reducing the data collection burden. For each sampled UE location in the training dataset, the groundtruth channel is generated by considering all 40 possible single-bounce paths via the scatterers. The delay for each path is calculated based on its total geometric length, which is the distance from the BS to the scatterer and then to the UE. The complex gain of each path is determined by two main factors: the path-length-dependent attenuation and a unique random reflection coefficient assigned to each scatterer. Although 40 paths are generated, the CADM is designed to learn the $L^{\prime}=5$ dominant ones. Therefore, a path selection mechanism is applied: the 5 paths with the highest power, as determined by their path gains, are chosen. The angles and delays of these 5 selected paths constitute the groundtruth data used for training the FCNN and the training loss is computed as the mean squared error (MSE) between the predicted and true channel parameters. The network was trained for 500 epochs using the Adam optimizer, with an initial learning rate of 0.01 and a batch size of 128.
@@
  • NLoS, Geometry-based: Assumes the shortest-delay path as LoS for geometric localization.
@@ VI Conclusion
This paper introduces a CKM-enabled ISAC framework that achieves robust environment-aware sensing, even in challenging NLoS scenarios where traditional geometry-based methods fail. We demonstrate that the same CKM designed for environment-aware communications can be exploited to achieve sensing as well. Simulation results demonstrate significant performance gains over geometry-based approaches, with CRLB analysis confirming enhanced robustness. This approach marks a significant advancement for ISAC systems, enabling seamless integration of sensing and communication in complex environments enabled by the same CKM.